OpenStudy (anonymous):

integral of ((1-x)/(x))^2 dx I mostly need help figuring out u and du.

zepdrix (zepdrix):

oh oh is the square on the whole thinggggg grrr wish i could read >.<

OpenStudy (anonymous):

yea it's on the whole thing

zepdrix (zepdrix):

\[\Large\rm \int\limits \left(\frac{1-x}{x}\right)^2dx\]Hmm I don't see a u-sub... Expand then simplify I guess?

zepdrix (zepdrix):

\[\Large\rm =\int\limits \frac{1}{x^2}-\frac{2}{x}+1~dx\]Something like that?

OpenStudy (usukidoll):

I think so and then you can integrate take antiderivative on each one

OpenStudy (usukidoll):

I see a logarithm

OpenStudy (anonymous):

Final answer I get is \[-\frac{ 1 }{ x } - 2\ln(x) +x + C\]

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (triciaal):